OpenStudy (haleyelizabeth2017):

The water level varies from 12 inches at low tide to 52 inches at high tide. Low tide occurs at 9:15 a.m. and high tide occurs at 3:30 p.m. What is a cosine function that models the variation in inches above and below the water level as a function of time in hours since 9:15 a.m. I have the first part of it, I just need helping finding the last bit of the function. So far I have 44 + 32 cos ______

OpenStudy (anonymous):

I don't think the 44 and 32 are right.

OpenStudy (haleyelizabeth2017):

Wait...20 is midline actually I think

OpenStudy (anonymous):

20 is amplitude

OpenStudy (haleyelizabeth2017):

Whoops...wasn't thinking lol 32 + 20 cos :P

OpenStudy (anonymous):

ok yeah

OpenStudy (haleyelizabeth2017):

:) how do I figure out the second part?

OpenStudy (anonymous):

the period is going to be twice the time between 9:15 am and 3:30 pm

OpenStudy (haleyelizabeth2017):

Which is 6.25 hours...so 13.5?

OpenStudy (haleyelizabeth2017):

And then I divide 2pi by that?

OpenStudy (anonymous):

12.5 hrs. and then yes divide 2π by that

OpenStudy (haleyelizabeth2017):

Ah! Sorry...brain's not wanting to work today :( did that earlier too haha so I got: \[32 + 20 cos \frac{\pi}{6.25}\] we didn't really even need to multiply it by 2 :P

OpenStudy (haleyelizabeth2017):

Wait! forgot x

OpenStudy (haleyelizabeth2017):

...or theta.

OpenStudy (anonymous):

yes. and I think there's a shift

OpenStudy (haleyelizabeth2017):

Ah. wouldn't that be x +/- something?

OpenStudy (haleyelizabeth2017):

Like it depends which way I thought...

OpenStudy (haleyelizabeth2017):

so \[32 + 20 cos \frac{\pi}{6.25}(x+9.25) \] maybe?

OpenStudy (anonymous):

yes. but since you shifted by half you could just switch the sign

OpenStudy (anonymous):

\[32-20\cos \frac{ \pi x}{ 6.25 }\]

OpenStudy (haleyelizabeth2017):

instead of adding 9.25? O.o

OpenStudy (anonymous):

yest

OpenStudy (haleyelizabeth2017):

Ah...is that all?

OpenStudy (anonymous):

yes

OpenStudy (haleyelizabeth2017):

Awesome! Thank you!

OpenStudy (anonymous):

you're welcome